substitution method:
x=y-1
y=4-2x?

Answer by ewatrrr(24785)


Hi
solving by substitution method:
x=y-1
y=4-2x
y = 4 - 2(y-1) |substituting for (y-1) for x
3y = 6
y = 2 and x = 1 (2-1)
CHECKING our Answer***
y = 4-2x
2 = 4 - 2